# RL-GSBridge: 3D Gaussian Splatting Based Real2Sim2Real Method for Robotic Manipulation Learning

### v1

@@ Abstract
Sim-to-Real refers to the process of transferring policies learned in simulation to the real world, which is crucial for achieving practical robotics applications. However, recent Sim2real methods either rely on a large amount of augmented data or large learning models, which is inefficient for specific tasks. In recent years, radiance field-based reconstruction methods, especially the emergence of 3D Gaussian Splatting, making it possible to reproduce realistic real-world scenarios. To this end, we propose a novel real-to-sim-to-real reinforcement learning framework, RL-GSBridge, which introduces a mesh-based 3D Gaussian Splatting method to realize zero-shot sim-to-real transfer for vision-based deep reinforcement learning. We improve the mesh-based 3D GS modeling method by using soft binding constraints, enhancing the rendering quality of mesh models. We then employ a GS editing approach to synchronize rendering with the physics simulator, reflecting the interactions of the physical robot more accurately. Through a series of sim-to-real robotic arm experiments, including grasping and pick-and-place tasks, we demonstrate that RL-GSBridge maintains a satisfactory success rate in real-world task completion during sim-to-real transfer. Furthermore, a series of rendering metrics and visualization results indicate that our proposed mesh-based 3D Gaussian reduces artifacts in unstructured objects, demonstrating more realistic rendering performance.
@@ I INTRODUCTION
Learning robotic action strategies in a simulated environment and then transferring them to real-world deployment represents an ideal robotic learning model, balancing both the cost and safety of the learning process. However, a significant bottleneck in this ideal model is the reliability of sim-to-real transfer, which impacts the potential and value of the entire framework and poses substantial challenges.
@@
With the continuous development of the sim2real field, extensive work is advancing the progress from multiple perspectives [1, 2, 3, 4]. However, most sim-to-real methods attempt to expand the distribution range of training data to cover various situations that may arise in reality or to train a highly generalized large-scale model to learn knowledge for different tasks. This significantly increases the difficulty of training, and in more complex tasks, the additional training burden often leads to learning failures.
@@
To avoid additional training burden and achieve ideal sim2real performance on specific tasks, a novel framework is needed. Recently, advances in radiance field-based reconstruction methods [5, 6, 7, 8] provide new directions for sim2real training. Based on a simple idea—using radiance field reconstruction to create a visually realistic robot training environment to achieve real2sim for vision-based policy training—can we achieve satisfactory sim2real performance? For this purpose, we design a new real2sim2real visual reinforcement learning framework, RL-GSBridge, that bridges the real-to-sim gap by 3D Gaussian splatting, as shown in Fig. 1. We utilize 3D Gaussian splatting and editing techniques, to provide a ‘virtual-reality’ simulation platform.
@@
Figure 1: Three Components of First-Person Perspective Manipulation in RL-GSBridge for Robotic Arms: (1) Real2Sim Transfer by soft mesh binding GS modeling. Given collected images from the real world, we utilize a novel soft mesh binding Gaussian models for high-quality image rendering while ensuring geometric accuracy. (2) Learn Policy at Simulator with GS Physics Dynamics-based GS Renderer. A physical dynamics-based GS editing and rendering method is proposed to enable RL algorithms to acquire realistic dynamic images. (3) Zero-shot Real-world Robot Manipulation. We directly apply the learned policy to real-world robotic manipulation tasks without any further training or fine-tuning.
@@
For vision-based robot tasks, it is crucial to avoid illusions caused by inconsistencies between visual perception and contact geometry. Thus, it is required to ensure an accurate geometric representation of the model while achieving more realistic rendering results. GaMeS [9] has drawn our attention as it is a mesh-based Gaussian Splatting (GS) method, which ensures that the optimization of the GS units is performed within the geometric mesh model. However, it enforces the binding of Gaussians to mesh grid planes in a manner that could be considered a hard mesh binding method, thereby constraining the flexibility of GS units. To address this problem, a soft mesh binding method is proposed for Gaussian Splatting to further enhance rendering performance while providing improved modeling and editing capabilities, even for non-rigid objects.
@@
Physics-based dynamics simulation is also a crucial and challenging aspect of sim2real. To tackle this, an off-the-shelf physics simulator is used to provide dynamic changing information. The GS editing process simultaneously updates the scene, ensuring that visual realism and the rendering results align with physical interaction processes.
@@
We use the designed simulator with the soft mesh binding based GS models to train a robotic arm on manipulation tasks with deep reinforcement learning (DRL). The tasks include object grasping and pick-and-place. These scenarios feature various textures, geometric shapes, and patterned desktop backgrounds, simulating a range of everyday desktop manipulation scenes.
@@
Under the RL-GSBridge framework, the policy trained in these environments demonstrates high sim2real few-shot transfer success rates. This means that the policy maintains effective performance on real-world tasks when deployed, reflecting a strong ability to generalize from simulation to actual environments.
@@ II-A Sim2Real Transfer in RL
RL constructs an interactive learning model where an agent learns to maximize expected rewards through trial and error. By incorporating deep learning, DRL enhances the framework’s ability to tackle more complex tasks [10]. DRL has shown impressive performance across various domains, including games [11], finance [12], autonomous driving [13], and robotics [14]. However, most applications are confined to virtual environments due to real-world constraints related to safety, efficiency, and cost.
To improve the feasibility of deploying models in the real world, many researchers strive to bridge the gap between simulation and reality[15]. Methods to bridge the gap include domain randomization [16] and domain adaptation [17]. Domain randomization involves varying task-related parameters in the simulation to cover a broad range of real-world conditions, while domain adaptation focuses on extracting a unified feature space from both sources. Higher-level learning methods include meta-learning [18] and distillation learning [19]. Meta-learning aims to teach robots the ability to learn new tasks, whereas distillation learning trains a student network using knowledge from expert networks.
Our approach aligns with the first category of gap-bridging methods, but with a unique twist. We use soft mesh binding GS to create highly realistic simulation environments for robot training. Typically, achieving such fidelity requires expensive 3D scanning equipment or CAD expertise. In contrast, GS models visually realistic simulation environments using only multi-view 2D images captured with consumer-grade devices.
@@ II-B Radiance Field in Robotics
Neural Radiance Fields (NeRF) [5] is a neural implicit representation technique for novel view synthesis. It optimizes the parameters through multi-view images, and allows for the synthesis of target views from any viewpoint using volumetric rendering. NeRF’s representation has been applied to various tasks, including Simultaneous Localization and Mapping (SLAM) [7], scene reconstruction [20], scene segmentation [21], navigation [22], and manipulation [23].
NeRF-RL [24] treats the novel view synthesis task of NeRF as a proxy task, where the learned encoding network is directly used as feature input for reinforcement learning. NeRF-dy [25] uses NeRF as a perceptual decoder for the hidden states of a world model, training an additional dynamics estimation network to predict future state changes. NeRF2Real [26] learns real-world contexts by converting background meshes into a simulator, and trains robots to perform visual navigation, obstacle avoidance, and ball-handling tasks. However, the training speed of Vanilla NeRF has consistently been a bottleneck limiting its further deployment in practical applications.
3D GS [27] is an explicit radiance field method that directly updates the attributes of each 3D Gaussian component to optimize scene representation. 3D GS employs a splatting technique for rendering, and achieve extremely fast training efficiency through CUDA parallel technology. Moreover, compared to the implicit representation of NeRF, the explicit representation facilitates tracking dynamic scene modeling and editing scene content.
ManiGaussian [28] uses 3D GS as a dynamic attribute encoder for the scene, encoding spatiotemporal dynamic features of the scene as input for reinforcement learning. Quach et al. [29] combine GS with Liquid networks for real-world drone flight navigation tasks. They train flight strategies using imitation learning in a simulator and then deploy them in the real world.
In contrast to NeRF2Real [26], utilizes the images from the MuJoCo simulator for texturing the foreground objects. We model each foreground object and use GS to edit the representation of the physical interactions with these foreground objects. Unlike the method designed by Quach et al. [29], which trains drone navigation for target-oriented tasks without involving complex interactions, our approach involves robotic arm manipulation tasks with complex interactions.
@@ III METHODS
The proposed framework aims to harness the potential of high-fidelity 3D GS models in sim2real for robot action training tasks. In this paper, we focus on manipulation tasks for robotic arms. As shown in Fig. 1, the overall framework is divided into two parts: real2sim and sim2real. In real2sim stage, we collect image data $\{{\mathbf{I}_{k}}\}$ in real-world operational scenarios, where $\mathbf{I}_{k}$ represent the image sequences of the $k$-th object or background in the scenarios. We will model both the geometry and appearance of the scene to build a simulation environment for the manipulation task. In sim2real stage, we use visual perception and deep reinforcement learning to train a policy network, and directly transfer the policy to the real world.
@@ III-A Real2Sim: Building simulator with soft mesh binding GS
To obtain a realistic training simulation environment, we use consumer-grade cameras to capture 2D image data $\mathbf{I}_{k}$ of desktop-level operating platforms. Our goal is to model a geometrically accurate and texture-realistic simulation model for training operational tasks. More specifically, we use mesh models $\{{\mathbf{M}_{k}}\}$ to represent the accurate geometric information, and Gaussian sets $\mathbf{G}_{k}=\{\mathbf{g}_{k}^{i}\}$ for high-quality texture. Below, we sequentially describe the steps to construct the Simulator with GS renderer.
@@ III-A1 Real-world Data Preparing
We use a monocular camera or mobile phone to capture a 1-2 minute video of the target object on the experimental platform. We select approximately 200 keyframes from the video and use COLMAP [30] to obtain the camera’s internal and external parameters for each frame. When modeling the object, image segmentation algorithm is also needed for extracting the target object. Many high-performance off-the-shelf segmenters are available, but we choose SAM-track [31] for efficient object segmentation.
@@ III-A2 Real2Sim modeling by soft mesh binding GS
here, $\mathbf{v}_{1},\mathbf{v}_{2},\mathbf{v}_{3}$ represents the triangular mesh vertex positions, $\alpha_{1}^{i},\alpha_{2}^{i},\alpha_{3}^{i}$ are learnable positional weight parameters for $\mathbf{g}_{i}$.However, we observe that this approach of enforced mesh-GS binding would diminish the flexibility of 3D Gaussians, limiting the optimization space of Gaussian units when the mesh model is not sufficiently fine or accurate, introducing certain undesirable defects.
To address this, we propose a soft mesh binding method, which builds on the rigid mesh binding of GaMeS [9] by relaxing the enforced hard binding and instead using a soft binding constraint to establish the association between mesh model elements and Gaussian units. We introduce a component along the normal direction into the mean vector of positions, specifically:
@@
here, $\alpha_{n}^{i}$ is a learnable weight parameter ranging in $[-1,1]$, and $\rm{\textbf{v}}_{n}$ is the normal vector. As shown in Fig. 2, our method allows the Gaussian units within the mesh to float within a certain range along the normal vector. This flexibility in Gaussian unit optimization could bring a smoother distribution of Gaussian units on the object’s surface. Ultimately, the algorithm not only ensures that the Gaussian model can still accurately represent the geometric structure of the scene but also offers some tolerance and refinement space, providing the possibility to handle non-rigid objects.
@@ III-A3 Physic Dynamics-Based GS Editing
We combine the dynamic simulation results from the simulator with real-time Gaussian model updates and render views to ensure that the visual representation follows the entire physical change process. We first use RANSAC plane regression and manual alignment methods to align GS models with mesh models in the simulator, and set the initial position of the GS model.
With the aligned initial model, we read the real-time pose changes of objects in the operational scene from the simulator. For the $k$-th object, We acquire the rotation quaternion $q_{k}$ and the homogeneous transformation matric $\mathbf{T}_{k}$ in the world coordinate system. Given Gaussian parameters $\theta_{k}^{i}=(\mu_{k}^{i},r_{k}^{i},s_{k}^{i},\sigma_{k}^{i},c_{k}^{i})$, The editing process of each Gaussian $\mathbf{g}_{k}^{i}$ in the Gaussian set $\mathbf{G}_{k}$ could be formulated as:
@@
After applying the correspondence transformations to all object models, we concat the Gaussian sets to acquire the Gaussian model of the whole scene $\mathbf{G}_{scene}$. Then we use rasterization to render $\mathbf{G}_{scene}$, obtaining the synchronized edited rendering view.
Figure 3: Manipulation Policy Training: Converting physical simulation dynamics signals into rotation and translation signals for GS editing, synchronizing GS model updates, and using first-person perspective scenes rendered by GS models as input for an actor-critic RL network to learn the manipulation policy.
TABLE I: Comparison of Success Rates for Small-cube and Bear Grasping Experiments between RL-GSBridge and RL-sim: Bold indicates the better result between the two methods, with values in parentheses representing the change in success rate when transferred to the real world relative to the simulation environment. (↓xx%) indicates a decrease of xx%, and (↑xx%) indicates an increase of xx%.
Object Small-cube Bear
@@
TABLE II: Sim2Real Results for the Grasping Task in Various Complex Manipulation Scenarios with RL-GSBridge.
Object (Background) Cake (Foam pad) Banana (Foam pad) Small-cube (Tablecloth) Cake (Tablecloth) Banana (Tablecloth) Bear (Tablecloth)
@@
We use Pybullet [32] as the simulation training platform and employ the SAC (Soft Actor-Critic) [33] algorithm for policy learning. The SAC is an offline policy algorithm belonging to maximum entropy RL, which consists of two critic networks, $Q_{\phi,1}(s,a)$ and $Q_{\phi,2}(s,a)$, and one actor network, $\pi_{\theta}(s)$. For the sampled set $B$ from the replay buffer, the update loss of critic is defined as:
@@
For continuous action spaces, the SAC policy network $\pi_{\theta}$ outputs actions following Gaussian distribution. The update loss of actor is defined as:
@@
Building on SAC and referencing DDPGwB [34], we propose SACwB, and introduce a lightly designed baseline controller incorporating artificial experience to guide the policy. This approach avoids complex reward designs while eliminating ineffective action spaces.
@@
In the supervision of the actor network, we introduce an additional behavior clone loss $L_{bc}$ in Equation 6 to supervise the mean of the actions output by the actor using base controller actions, with the probability $\lambda$:
@@
The whole training pipeline in the simulator with physic dynamics-based GS renderer is shown in Fig. 3. The RL algorithm provides executable actions to the physics simulator, which will provide state information for actor and critic learning. The GS renderer simultaneously renders the realistic images by physic dynamics-based GS rendering, serving as observation input to the actor network. We rely solely on the first-person perspective 2D image and proprioceptive state of the robotic arm, considering several advantages as described in [35]. To enhance sim2real robustness, we add random noise to the rendered images of the GS model and randomly alter certain image attribute parameters during training for environment challenges.
After policy training in the simulator, we use ROS as the communication medium, with the actual first-person perspective camera observations of the environment serving as input to the policy inference network. The trained policy network then outputs executable actions.

### v2

@@ Abstract
Sim-to-Real refers to the process of transferring policies learned in simulation to the real world, which is crucial for achieving practical robotics applications. However, recent Sim2real methods either rely on a large amount of augmented data or large learning models, which is inefficient for specific tasks. In recent years, with the emergence of radiance field reconstruction methods, especially 3D Gaussian splatting, it has become possible to construct realistic real-world scenes. To this end, we propose RL-GSBridge, a novel real-to-sim-to-real framework which incorporates 3D Gaussian Splatting into the conventional RL simulation pipeline, enabling zero-shot sim-to-real transfer for vision-based deep reinforcement learning. We introduce a mesh-based 3D GS method with soft binding constraints, enhancing the rendering quality of mesh models. Then utilizing a GS editing approach to synchronize the rendering with the physics simulator, RL-GSBridge could reflect the visual interactions of the physical robot accurately. Through a series of sim-to-real experiments, including grasping and pick-and-place tasks, we demonstrate that RL-GSBridge maintains a satisfactory success rate in real-world task completion during sim-to-real transfer. Furthermore, a series of rendering metrics and visualization results indicate that our proposed mesh-based 3D GS reduces artifacts in unstructured objects, demonstrating more realistic rendering performance.
@@ I INTRODUCTION
Learning robotic action policies in simulation and transferring them to real-world represents an ideal robotic learning strategy that balances both the cost and safety of the learning process. However, a significant bottleneck is the reliability of sim-to-real transfer, which impacts the potential of the entire framework towards substantial challenges.
@@
With the continuous development of the simulation-to-reality(Sim2Real) field, extensive work is advancing progress from multiple perspectives [1, 2, 3, 4]. However, most Sim2Real methods attempt to expand the distribution of training data to cover various situations that may arise in reality, or to train a highly generalized large model to learn knowledge for different tasks. This significantly increases the difficulty in training stage.
@@
To avoid additional training burden and achieve ideal Sim2Real performance on specific tasks, a novel framework is needed. Recently, advances in radiance field-based reconstruction methods [5, 6, 7, 8] provide new directions for Sim2Real training. Based on a simple idea—using radiance field reconstruction to create a visually realistic robot training environment—can we achieve satisfactory Sim2Real performance? For this purpose, we design a Real2Sim2Real visual reinforcement learning framework, RL-GSBridge, that bridges the real-to-sim gap by 3D Gaussian splatting, as shown in Fig. 1. Utilizing 3D Gaussian splatting and editing techniques, RL-GSBridge provide a ‘virtual-reality’ simulation platform for policy learning.
@@
Figure 1: Pipeline of RL-GSBridge. (1) Real2Sim Environment Transfer. Real-world scenarios is reconstructed through a novel soft mesh binding GS model. (2) Learn Policy at Simulator with GS Render. With physical dynamics-based GS editing, RL policies learn through realistic rendered images in simulation. (3) Zero-shot Real-world Robot Manipulation. We directly apply the policy to real-world tasks without fine-tuning.
@@
For vision-based robot tasks, it is crucial to avoid illusions caused by inconsistencies between visual perception and contact geometry. Thus, it is required to ensure an accurate geometric representation of the model while achieving more realistic rendering results. GaMeS [9] has drawn our attention as it is a mesh-based Gaussian splatting (GS) method, which ensures that the optimization of GS units is performed within the geometric mesh model. However, it enforces Gaussians to be aligned with the mesh grid planes, which could be considered as ‘hard mesh binding’, thereby limiting the flexibility of GS units. To address this, we propose a soft mesh binding method for Gaussian Splatting, which could further enhance rendering quality while preserving editing capabilities for both objects and the background.
@@
Physics-based dynamics simulation is also a crucial and challenging aspect of sim2real. To tackle this, an off-the-shelf physics simulator is used to provide dynamic changing information. The GS editing process simultaneously updates the scene, ensuring that the rendering results align with physical interaction processes.
@@
Based on the designed simulation platform, we train robotic manipulation policies by deep reinforcement learning methods. The model is trained on grasping and pick-and-place tasks across scenarios that include diverse textures, geometric shapes, and patterned desktop backgrounds.
@@
Under the RL-GSBridge framework, the policy shows a minor variation in success rates during Sim2Real transfer. This means that the policy maintains effective performance on real-world tasks, reflecting a strong ability to generalize from simulation to real-world environments.
@@ II-A Sim2Real Transfer in RL
RL constructs an interactive learning model in which an agent learns to maximize rewards through trial and error. By incorporating deep learning, deep reinforcement learning(DRL) enhances the framework’s ability to tackle more complex tasks [10]. DRL has shown impressive performance across various domains, including games [11], finance [12], autonomous driving [13], and robotics [14]. However, most applications are restricted to virtual environments due to real-world constraints related to safety, efficiency, and cost.
To improve the feasibility of deploying models in the real world, many researchers strive to bridge the gap between simulation and reality[15]. These methods include domain randomization [16] and domain adaptation [17, 18]. Domain randomization involves varying task-related parameters in the simulation to cover a broad range of real-world conditions, while domain adaptation focuses on extracting a unified feature space from both sources. Higher-level learning methods include metalearning [19] and distillation learning [20]. Meta-learning aims to teach robots the ability to learn new tasks, whereas distillation learning trains a student network using knowledge from expert networks.
Our approach aligns with the first category of gap-bridging methods, but with a unique twist. We use soft mesh binding GS to create realistic simulation environments for robot training. Typically, achieving such fidelity requires expensive 3D scanning equipment or CAD expertise. In contrast, GS models visually realistic simulation environments using only multi-view images captured with consumer-grade devices.
@@ II-B Radiance Field in Robotics
Neural Radiance Fields (NeRF) [5] is an implicit representation technique for novel view synthesis. It optimizes the parameters through multi-view images, and allows for the synthesis of any target views using volumetric rendering. NeRF’s representation has been applied to various tasks, including Simultaneous Localization and Mapping (SLAM) [7, 21, 22], scene reconstruction [23], scene segmentation [24], navigation  [25], and manipulation [26].
NeRF-RL [27] treats the novel view synthesis task of NeRF as a proxy task, where the learned encoding network is directly used as feature input for reinforcement learning. Y. Li et al. [28] uses NeRF as a perceptual decoder for the hidden states of a world model, training an additional dynamics estimation network to predict future state changes. NeRF2Real [29] learns real-world contexts by converting background meshes into a simulator, and trains robots to perform visual navigation, obstacle avoidance, and ball-handling tasks. However, the training and rendering speed of Vanilla NeRF has consistently been a bottleneck limiting its further deployment in practical applications.
3D GS [30] is an explicit radiance field method that directly updates the attributes of each 3D Gaussian component to optimize scene representation. 3D GS employs a splatting technique. for rendering, and achieve extremely fast training efficiency through CUDA parallel technology. Moreover, compared to the implicit representation of NeRF, the explicit representation facilitates tracking dynamic scene modeling and editing scene content.
Many robotics works also incorporate 3D Gaussian techniques for perception and motion learning [31]. ManiGaussian [32] uses 3D GS as visual and dynamic scene representation for policy learning. Quach et al. [33] combine GS with Liquid networks for real-world drone flight navigation tasks, training in simulation and then deploying the policy in the real world.
In contrast to NeRF2Real [29], which directly textures the foreground objects in simulator, we model each foreground object with editable GS parameters. Also unlike the method designed by Quach et al. [33], which trains drone navigation policies for target-oriented tasks, our approach involves robotic arm manipulation tasks with complex interactions.
@@ III METHODS
RL-GSBridge aims to harness the potential of high-fidelity 3D GS models in Sim2Real for robot action training tasks. In this paper, we focus on manipulation tasks for robotic arms. As shown in Fig. 1, the overall framework is divided into two parts: Real2Sim and Sim2Real. In Real2Sim stage, we collect real-world image data $\{{\mathbf{I}_{k}}\}$, where $\mathbf{I}_{k}$ represent the image sequences of the $k$-th object or background in the scenarios. We will model both the geometry and appearance of the scene to build a simulation environment for the manipulation task. In Sim2Real stage, we use visual perception and deep reinforcement learning to train a policy network, and directly transfer the policy to the real world.
@@ III-A Real2Sim: Building simulator with soft mesh binding GS
To obtain a realistic simulation environment, we use consumer-grade cameras to capture 2D image data $\mathbf{I}_{k}$ of desktop-level operating platforms. Our goal is to model a geometrically accurate and texture-realistic simulation model for training operational tasks. More specifically, we use mesh models $\{{\mathbf{M}_{k}}\}$ to represent the accurate geometric information, and Gaussian sets $\mathbf{G}_{k}=\{\mathbf{g}_{k}^{i}\}$ for high-quality texture. Below, we sequentially describe the steps to construct the Simulator with GS renderer.
@@ III-A1 Real-world Data Preparing
We use a monocular camera or mobile phone to capture a 1-2 minute video of the target object on the experimental platform. We select approximately 200 keyframes from the video and use COLMAP [34] to obtain the camera’s internal and external parameters for each frame. Image segmentation algorithm is also needed for extracting the target object, and we use an off-the-shelf segmenter, SAM-track [35], for efficient object segmentation.
@@ III-A2 Real2Sim modeling by soft mesh binding GS
here, $\mathbf{v}_{1},\mathbf{v}_{2},\mathbf{v}_{3}$ represents the triangular mesh vertex positions, $\alpha_{1}^{i},\alpha_{2}^{i},\alpha_{3}^{i}$ are learnable positional weight parameters for $\mathbf{g}_{i}$. However, this approach of enforced mesh-GS binding would diminish the flexibility of 3D Gaussians, limiting the optimization of Gaussian units when the mesh model is not accurate, introducing certain undesirable defects.
To address this, we propose a soft mesh binding method, which builds upon GaMeS [9], but relaxes the enforced hard mesh binding to a soft binding constraint. We introduce a component along the normal direction into the vector of positions, specifically:
@@
here, $\alpha_{n}^{i}$ is a learnable weight parameter and $\rm{\textbf{v}}_{n}$ is the normal vector. $\alpha_{n}^{i}$ is constrained within the range of $[-1,1]$, ensuring the association between each mesh model element and Gaussian pairs. As shown in Fig. 2, our method allows the Gaussian units within the mesh to float within a certain range along the normal vector. This flexibility in Gaussian unit optimization could bring a smoother distribution of Gaussian units on the object’s surface. Ultimately, the algorithm not only ensures that the Gaussian model can still represent the accurate geometric structure according to the mesh models, but also offers some tolerance and refinement space. Besides, the binds between the mesh grid and Gaussians could even provide the possibility to handle non-rigid objects, as demonstrated in section IV-B4.
@@ III-A3 Physic Dynamics-Based GS Editing
After obtaining the visual GS models $\{\mathbf{G}_{k}\}$ and geometry models $\{\mathbf{M}_{k}\}$ for real-world objects, we combine the dynamic simulation results from the simulator with real-time Gaussian model updates and render views to ensure that the visual representation follows the entire physical change process. We first use RANSAC plane regression and manual alignment methods to align GS models with mesh models in the simulator, and set the initial position of the GS model.
With the aligned initial model, we read the real-time pose changes of each object in the operational scene from the simulator. For the $k$-th object with its GS model $\{\mathbf{G}_{k}\}$, We acquire the rotation quaternion $q_{k}$ and the homogeneous transformation matric $\mathbf{T}_{k}$ in the world coordinate system. Given Gaussian parameters $\theta_{k}^{i}=(\mu_{k}^{i},r_{k}^{i},s_{k}^{i},\sigma_{k}^{i},c_{k}^{i})$, The editing process of each Gaussian $\mathbf{g}_{k}^{i}$ in the model set $\mathbf{G}_{k}$ could be formulated as:
@@
As for local non-rigid transformations on the mesh grid, the Gaussians could be updated according to Equation 2. After applying transformations to all object models, we concat the Gaussian sets to acquire the Gaussian model of the whole scene $\mathbf{G}_{scene}$. Then we use rasterization to render $\mathbf{G}_{scene}$, obtaining the synchronized edited rendering view.
Figure 3: Policy training pipeline in RL-GSBridge. In the upper half of the figure, physic dynamics-based GS editing receives the transformation signals of objects and synchronizes the states of GS models. In the lower half of the figure, an actor-critic RL network receives first-person perspective images rendered by GS models as input, to learn a vision-based manipulation policy.
TABLE I: Comparison of Success Rates Between RL-GSBridge and RL-Sim in Grasping Experiments, All Conducting Under Foam Pad(FP) Background. Bold Indicates Better Results. Values in Parentheses Represent Relative Change in Success Rate During Sim2Real Transfer. (↓xx%) Indicates a Decrease, While (↑xx%) Indicates an Increase.
Object Small_cube Bear
@@
TABLE II: Sim2Real Results for Grasping Task in Various Manipulation Scenarios. Contents in Parentheses After the Object Names Represent Different Backgrounds(Bg), Where FP Denotes Foam Pad, and TC Denotes Tablecloth.
Object (Bg) Cake (FP) Banana (FP) Small_cube (TC) Cake (TC) Banana (TC) Bear (TC)
@@
We use Pybullet [36] as the simulation training platform, and employ SAC (Soft Actor-Critic) [37] algorithm for policy learning, due to its mature development in RL and widespread application in robotics. The SAC is an offline policy algorithm belonging to maximum entropy RL, which consists of two critic networks, $Q_{\phi,1}(s,a)$ and $Q_{\phi,2}(s,a)$, and one actor network, $\pi_{\theta}(s)$. For the sampled set $B$ from the replay buffer, the update loss of critic is defined as:
@@
In continuous action space, policy $\pi_{\theta}$ outputs actions of Gaussian distribution. The loss for actor is defined as:
@@
Since vanilla SAC struggles to converge rapidly under sparse reward, to accelerate the learning process through automated guidance, we propose SACwB by referring to DDPGwB [38], and introduce a lightly designed baseline controller to guide the policy. This approach avoids complex reward designs while eliminating ineffective action spaces.
@@
For the supervision of the actor network, we introduce an additional behavior clone loss $L_{bc}$ in Equation 6 to supervise the mean values of the action distribution through base controller actions, with the probability $\lambda$:
@@
The whole training pipeline in the simulator with physic dynamics-based GS renderer is shown in Fig. 3. The RL algorithm provides executable actions to the physics simulator, which will return state information for actor-critic learning. The GS renderer simultaneously renders the realistic images by physic dynamics-based GS rendering, serving as observation input to the actor network. We rely solely on the first-person perspective image and proprioceptive state of the robotic arm, considering several advantages as described in [39]. To enhance sim2real robustness, we add random noise to the rendered images of the GS model and randomly alter certain image attribute parameters during training for environment challenges.
After policy training in the simulator, we directly deploy the actor network onto the real robot arm, with the real-world eye-in-hand camera observations of the environment serving as visual input. The trained policy subsequently outputs the position of the end-effector and the gripper states as executable actions.